Database last updated on Mon, 02 Feb 2026 03:40:49 +0000 by localhost
Instrument constant changes for the week before 2023265 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 18 rows)
Measurements and tests summary.
| _jday | _loc | _serial | _skc |
| 265 | TORONTO | #191 | calds |
(0 sec)
Dead time (0 sec)
Nothing to report.
Standard lamp R6 (1 sec)
Nothing to report.
Standard lamp R5 (1 sec)
Nothing to report.
Number of days since last dome/window cleaning (1 sec, 2 rows)
| serial | days |
| 191 | more than a year |
Run/Stop results for last week. (0 sec, 1 rows)
| Serial | Total | Failed 0 | Failed dark | Failed 1 | Failed 2 | Failed 3 | Failed 4 | Failed 5 | Failed double |
| 191 | 12 | 0 | 0 | 0 | 1 | 1 | 0 | 0 | 0 |
Ozone results for good DS obs (std<2.5DU, mu<4) within 3 minutes between 2 Brewers. (0 sec)
No ozone measurements to display.
Unsorted comments (1 sec, 10678 rows)
| Serial | type | Text | Count | YYYYJJJ |
| 191 | env: | DosBox | 1 | 2023265 |
| 191 | cubdsp: | 1 | 1 | 2023265 |
| 191 | fre_after: | checkaz 7717 7739 | 1 | 2023265 |
| 191 | fre_after: | po 2114 2135 | 1 | 2023265 |
| 191 | fre_after: | re 3168 3179 | 1 | 2023265 |
| 191 | re: | BREWER SPECTROPHOTOMETER #00191 AES SCI | 1 | 2023265 |
The END (0 sec, 2 rows)